OpenStudy (anonymous):

solve inseries d2y/dx2+x2y=0

OpenStudy (dumbcow):

separate variables \[\frac{ dy^2}{y} = - x^2 dx^2\] integrate twice \[(\ln y +m) dy = (-\frac{x^3}{3} + k) dx\] \[y \ln y - y + my = -\frac{x^4}{12} +kx +C\] \[y = \frac{C +kx -x^4}{12( \ln y - 1+m)}\] because of the ln(y) you cant algebraically solve for 'y' explicitly using elementary functions